4. ) Calculate the density of H2S gas at 56 oC and 967 mmHg. Obtain the density in grams per liter.
5. A 2.30g sample of a white solid was vaporized in 345 mL vessel. If the vapor has a pressure of 985 torr at 148 oC, what is the molar mass of the solid?
6. A 3.75 L sample of gas from a chemical plant is collected and analyzed at 25oC. It is found to contain 0.0674 g NO, 0.00459 g C2H6, 0.00275 g CO2, and 0.185 g of H2O.
a. What are the partial pressures in atmospheres of each component of gas?
b. What is the total pressure in mmHg of the sample?
c. What is the mole fraction of each gas?
1
Expert's answer
2019-12-12T05:46:27-0500

4.Assume you have 1 mole of the gas and calculate the new volume under the temperature and pressure given.

P1*V1/T1 = P2*V2/T2

Solve for V2: P1*V1*T2/T1*P2 = V2

Convert the units as necessary:

56 C = 273 + 56 K = 329 K

967 mmHg = 967/760 atm = 1.272 atm

Now, plug the numbers into the formula:

1atm x 22.4L x 329K/273K x 1.272atm = 21.2 liters

Now, we have one mole of gas occupying 21.2 liters, so we can easily determine the density:

H2S = (2 x 1.01) + 32.06g/mole = 34.08g/mole

(34.08g/mole)/(21.2L/mole) = 1.61 g/L

5.177 g/mol

6.We know that in a gaseous mixture, the partial pressure exerted by a component gas is the same as the pressure it would exert if it alone occupied the container. The total pressure is the sum of the individual partial pressures.